please graph this logarithemic function.. thanks big time:)
[] means the number in it is the base..

1. y=3-log[3]x

Answer by anantha(86)

sol:
y=3-log[3]x
lo[3]x=3-y
x=3^(3-y) (i.e;log[a]x=b then x=a^b)
here we are taking the values for y
let y=-1,then the value of x=3^(3-(-1))
x=3^(3+1)
x=3^4=81
(x,y)=(81,-1)
let y=0,then the value of x=3^(3-0)=3^3=27
(x,y)=(27,0)
let y=1 then the value of x=3^(3-1)=3^2=9
(x,y)=(9,1)
let y=2 then the value of x=3^(3-2)=3^1=3
(x,y)=(3,2)
let y=3 then the value of x=3^(3-3)=3^0=1
(x,y)=(1,3)
the points are(81,-1),(27,0),(9,1),(3,2),(1,3)
plotting the above points we get the requried graph